OpenStudy (anonymous):

Find the values of m which one root of the equation 4x^2=mx-5 is three times the other root.

OpenStudy (lalaly):

in a quadratic equation \[ax^2+bx+c=0\] sum of roots=-b/a product of roots=c/a here a=4 b=-m c=5 let the roots be L and N L=3N so now tha roots are N and 3N sum of roots=>> N+3N=m/4 4N=m/4 N=m/16 Product of roots =>> N*3N=5/4 3N^2=5/4 Substitute the first equation in the second 3(m/16)^2=5/4 and solve for m

OpenStudy (anonymous):

Thank you

OpenStudy (lalaly):

=)

hero (hero):

Impressive lalaly. You're the smartest girl on here besides myininaya

OpenStudy (lalaly):

i know im not, But thanks hero :$

hero (hero):

don't doubt yourself

OpenStudy (lalaly):

Thats sweet, Merci :$